# CHARME: A chain-based reinforcement learning approach for the minor embedding problem

### v1

@@ 4.1. Setup
Dataset. We conduct our experiments on synthetic and real datasets. For the synthetic dataset, we generate logical graphs using Barabási-Albert (BA) model. In details, we split the synthetic dataset into a training set and a testing set. The training set, denoted as $\mathcal{G}_{syn}$, includes 700 BA graphs with $n=150$ nodes and degree $d=10$, while the testing set includes 300 new graphs which are generated by varying the number of nodes $n\in\{80,100,120,140\}$ and the degree $d\in\{2,5,10\}$. The hardware graph used in this experiment is in form of Chimera topology which is a $45\times 45$ grid of $K_{4,4}$ unit cells (Boothby et al., 2016). The experiment on the synthetic dataset aims to evaluate the performance of CHARME on logical graphs from the same distribution.
Furthermore, we also conduct the experiments on the real dataset which is generated based on the QUBO formulations of test instances of the Target Identification by Enzymes (TIE) problem proposed in the work (Ngo et al., 2023b). In particular, different metabolic networks for the TIE problem result in different QUBO formulations. As we mentioned in previous sections, each specific QUBO formulation corresponds to a logical graph. Thus, we can generate various logical graphs by varying the metabolic networks for the TIE problem. In this experiment, we collect 105 metabolic networks from the KEGG database (Kanehisa and Goto, 2000). Then, we generate corresponding logical graphs by using the QUBO formulations with these metabolic networks as input. In collected graphs, the number of nodes ranges from 6 to 165, while the number of edges ranges from 22 to 578. We split the dataset into three groups based on the ratio between the number of nodes $n$, and the number of edges $m$: low density group ($m\leq 2n$), medium density group ($2n<m\leq 3n$), and high density group ($m>3n$). We use 70 graphs for training, and preserve the remaining for testing. We denote the training set $\mathcal{G}_{real}$. This experiment aimed to illustrate the effectiveness of our hybrid RL model in real-world scenarios.
Benchmark. To demonstrate the effectiveness of our Order Exploration strategy, we establish a comparative greedy strategy referred to as Greedy Exploration. Greedy Exploration approach explores the embedding orders based on the $B^{*}$ algorithm.
@@
  • OCT-based (Goodrich et al., 2018): This method follows the top-down approach that minimizes the number of qubits based on their proposed concept of the virtual hardware framework. In this experiment, we use the most efficient version of OCT-based, named Fast-OCT-reduce. While the OCT-based is effective in reducing the number of qubits, it suffers from high computational complexity, leading to long running time to find a solution. This can pose a bottleneck for quantum annealing process, especially when frequency of quantum annealing usage has been increased rapidly.
  • Minorminer (Cai et al., 2014): This method is a heuristic method developed commercially by D-Wave. This method finds solutions by adding nodes of a logical graph to the hardware graph sequentially. Minorminer is an efficient approach for sparse logical graphs.
@@
  • ATOM (Ngo et al., 2023a): This method is a heuristic method which finds solutions based on the concept of adaptive topology. The strength of this method is the ability to find solutions in small running time without compromising their quality. The concept of adaptive topology is integrated in CHARME through the state transition algorithm. Thus, by comparing its performance to that of ATOM, we can assess the effectiveness of RL in leveraging a heuristic method.
@@
Hyper-parameter selection. In the exploration phase, we apply Exploration Order, described in Algorithm 2, to precompute embedding orders for each graph $P_{i}\in\mathcal{G}_{syn}\cup\mathcal{G}_{real}$. Specifically, we set the sampling $D$ as $10000$, the exploration limit $K$ as $10^{6}$ and use up to $1024$ threads running in parallel for exploring embedding orders.
@@
For training, the method for GNN layers in both actor and critic models is graph convolutional networks (Kipf and Welling, 2017). Each model has three GNN layers. The learning rate for the actor model is $3\times 10^{-4}$, while the learning rate for the critic model is $10^{-5}$. The number of episodes to update policy is $100$, and the batch size in each update is $290$.
@@
For inference, due to the stochastic in behaviours of machine learning models, we sample $10$ solutions for each testcase, and report the solution with the least number of qubits. In addition, we use multi-processing for sampling solutions, so the actual running time of inference is equal to the time to sample one solution.
Figure 5. The figure highlights the performance of Order Exploration and Greedy Exploration in discovering embedding orders for training graphs in synthetic training set under various settings of $(n,d)$. In subfigures, the x-axis denotes the number of exploration steps, while the y-axis represents the corresponding efficiency scores of the resulting orders. As a result, in these subfigures, each orange (blue) point indicates the best efficiency score of the embedding order resulted by Order Exploration (Random Exploration) after a specific number of exploration steps.
Figure 6. The figure highlights the performance of Order Exploration and Greedy Exploration ($B^{*}$ style) in discovering embedding orders for training graphs in the real training set under various graph types including low density, medium density and high density. In subfigures, the x-axis denotes the number of exploration steps, while the y-axis represents the corresponding efficiency scores of the resulting orders. As a result, in these subfigures, each orange (blue) point indicates the best efficiency score of the embedding order resulted by Order Exploration (Random Exploration) after a specific number of exploration steps.
4.2. Results
4.2.1. The performance of the proposed exploration strategy
This section demonstrates the utility of Order Exploration in generating precomputed actions, commonly referred to as embedding orders, to streamline the training process of CHARME. Our experiments are designed to assess two key aspects: 1) the effectiveness of Theorem 1 with general prefix that implies the the practical significance of the lower bound outlined in Equation 15, and 2) the overall performance of our proposed exploration strategy in terms of the efficiency scores of resulting embedding orders.
@@ 4.2.1. The performance of the proposed exploration strategy
The effectiveness of Theorem 1 with general prefix. In this experiment, we designed a meticulous procedure to evaluate the effectiveness of Theorem 1 for general prefix orders which are not necessarily non-expansion. Specifically, we compare the number of qubits gained when embedding a same node $v$ at two different embedding steps $t$ and $t+\delta$, denoted as $C(v|t)$ and $C(v|t+\delta)$. In Theorem 1, we prove that $C(v|t)\leq C(v|t+\delta)$ if there is no expansion embedding between $t$ and $t+\delta$. Thus, our focus here lies on comparing $C(v|t)$ and $C(v|t+\delta)$ when at least one expansion embedding exists between these steps. Given a graph $P=(V_{P},E_{P})\in\mathcal{G}_{syn}\cup\mathcal{G}_{real}$, the procedure for selecting $v$, $t$, and $\delta$ unfolds as follows. First, we randomly select a subset $S_{P}\subset V_{P}$ following by selecting a random prefix $O_{A}\in\mathcal{E}^{(S_{P})}$. We select $v$ as a random node in $V_{P}\setminus S_{P}$. Next, we need to select $t$ and $\delta$ such that there exists one expansion embedding from $t$ to $t+\delta$. In details, we start with iteratively embedding nodes following the prefix order $O_{A}$. Subsequently, we randomly embed nodes $v^{\prime}\in V_{P}\setminus S_{P},v^{\prime}\neq v$ until we obtain an expansion embedding at step $|S_{P}|+\tau$. Finally, $t$ is assigned as $|S_{P}|+\tau-1$ and $\delta$ is randomly selected such that $\delta\geq 1$. By repeatedly performing this procedure, we can obtain different $v$, $t$, and $\delta$ for each training graph. For each triplet $(v,t,\delta)$ of a given training graph, $C(v|t)$ and $C(v|t+\delta)$ is calculated similarly as discussed in Theorem 1. For each setting of $(n,d)$, we measure the average of $C(v|t)$ and $C(v|t+\delta)$ on all triplet $(v,t,\delta)$ sampled from graphs with the setting $(n,d)$. We denote the average on $C(v|t)$ and $C(v|t+\delta)$ as early embedded gain and late embedded gain respectively.
@@
The results are illustrated in Table 2. Across all $(n,d)$ configurations, the average qubit requirement for embedding a node prior to topology expansion (early embedded gain) is consistently lower compared to post-topology expansion (late embedded gain). Interestingly, when the graph’s density increases, the difference in qubit gain between before and after topology expansion become more significant. This trend can be attributed to the expansion of the topology, which, on average, elongates the length of clean paths. Consequently, embedding a node after topology expansion tends to incur higher costs. These observations highlight the significance of the lower bound $\bar{F}$ on the efficiency score of embedding orders with a general prefix, as derived from Equation 15.
@@
The efficiency in exploring orders. In this experiment, we conducted a comparison between Order Exploration and Greedy Exploration to assess their effectiveness in exploring embedding orders for training logical graphs across various node sizes and degrees $(n,d)$. Figure 5 presents subfigures corresponding to each setting $(n,d)$ of logical graphs in the synthetic training set $\mathcal{G}_{syn}$. Each subfigure showcases the efficiency scores of embedding orders resulted from both strategies with variations in the exploration limit $K$ which is mentioned in Algorithm 2. We recall that given a logical graph $P$ and a hardware graph $H$, the efficiency score of an embedding order $O$ is the number of qubits when embedding $P$ into $H$ following the order $O$.
@@
The results depicted in the subfigures highlight the superior performance of Order Exploration over Greedy Exploration across all $(n,d)$ settings. In detals, the disparity becomes more significant with increasing values of $n$ and $d$. For example, with $(n,d)=(80,2)$, the average efficiency score of Order Exploration is approximately $2\%$ less than that of Random Exploration, while with $(n,d)=(140,10)$, the gap increases to around $9\%$. Furthermore, we observe that the disparity also widens as the exploration limit $K$ increases from 00 to $10^{6}$. The key difference between two methods is the lower bound derived from Theorem 1. Thus, by these results, we assert the practical utility of Theorem 1 in guiding the exploration of embedding orders, particularly in scenarios involving large $(n,d)$ values and high exploration limits $K$.
@@
Similarly, when applied to the real training set $\mathcal{G}_{real}$, the Order Exploration approach consistently outperforms Greedy Exploration in terms of efficiency scores (Figure 6). Specifically, when dealing with low-density and medium-density graphs, we observe that the disparity between Order Exploration and Greedy Exploration is small. This is because the small number of nodes and edges limit the available options for embedding orders. However, when dealing with high-density graphs, Order Exploration significantly outperforms Greedy Exploration. Specifically, Order Exploration requires approximately $1100$ qubits, while Greedy Exploration requires more than $1600$ qubits for its solution. This illustrates the superior performance of Order Exploration, particularly in settings involving high-density graphs.
@@ 4.2.2. The performance of CHARME
In this experiment, we compare the performance of CHARME  against three state-of-the-art methods in the synthetic and real datasets.
n 80 100 120 140 degree 2 5 10 2 5 10 2 5 10 2 5 10 OCT-based 357.76 845.38 1222.11 515.74 1263.32 1850.9 725.84 1782.18 2607.06 931.24 2348.74 3488.18 Minorminer 338.56 1236.88 2110.16 501.56 1914.18 3313.24 664.76 2632.82 4854.10 830.56 3554.01 6573.84 ATOM 384.06 1126.14 1815.72 546.62 1782.86 2882.72 727.46 2560.36 4217.48 998.58 3448.24 5621.88 CHARME 324.62 1027.36 1686.20 485.16 1580.08 2566.14 655.88 2263.66 3870.02 821.88 3082.59 5100.54
@@
Table 3. The average qubit usage of four methods including OCT-based, Minorminer, ATOM, and CHARME  corresponding to 12 different graphs’ sizes in the synthetic dataset.
@@
Figure 7. Comparison between four methods in term of running time.
@@
The synthetic dataset. Here, we assess the performance of CHARME  approach in the synthetic dataset, focusing on the number of qubits used and the running time. The goals of this experiment are as follows:
@@
  • We evaluate how effectively CHARME  framework leverages the heuristic method ATOM.
@@
  • We compare the quality of solutions in terms of the number of qubits obtained by CHARME  approach with three state-of-the-art methods.
@@
  • We measure the inference time of CHARME  approach and compare it with the running time of other methods to assess its computational efficiency.
@@
Table 3 presents the average number of qubits obtained from solutions generated by four methods for various pairs of size and degree $(n,d)$ of logical graphs. First of all, the results imply that CHARME  leverages the heuristic strategy used in ATOM. Across all pairs $(n,d)$, CHARME consistently provides solutions with fewer qubits compared to ATOM. In particular, CHARME achieves 11.36% reduction in qubit usage compared to ATOM. This observation not only indicates the efficiency of the proposed CHARME framework, but it also illustrates the generalization ability of the RL framework because the policy learnt from the training set can perform well for new graphs with different sizes and degrees.
@@
Turning to solution quality, the results in Table 3 demonstrate an out-performance three state-of-the-art methods in term of qubit usage in test cases with degree $d=2$. On average, in these cases, CHARME achieves 2.52% of reduction in term of qubit usage, compared to current best methods. In other cases, although CHARME cannot overcome the OCT-based in term of qubit usage, it still outperforms Minorminer and ATOM. This observation, combined with the subsequent analyse of running time is important to illustrate the practical effectiveness of CHARME in handling the minor embedding problem.
@@
Finally, we examine the running time of four methods. The results in Figure 7 illustrate that the running time of OCT-based is significantly larger than the running time of three other methods. In particular, the running time of OCT-based is approximately 6 times bigger than the running time of the second slowest method, Minorminer. Thus, although OCT-based is efficient in term of qubit usage in some certain cases, its expensive running time can pose a significant bottleneck in practical quantum annealing process. On the other hand, the running time of CHARME is comparable to the fastest method, ATOM. Furthermore, as we mentioned above, CHARME can offer better solutions than Minorminer and ATOM in all cases. Therefore, CHARME can be the optimal choice for tackling minor embedding problem in practice, especially when current quantum annealers must handle user requests immediately due to the raising demand on their utilization.

### v2

@@ 4.1. Setup
Dataset. We conduct our experiments on synthetic and real datasets. For the synthetic dataset, we generate logical graphs using the Barabási-Albert (BA) model. We then categorize the logical graphs of the synthetic dataset into two independent sets: a training set and a test set. The training set, denoted as $\mathcal{G}_{syn}$, includes $70$ BA graphs with $n=150$ nodes and degree $d=10$. The test set consists of $600$ new graphs that are generated by varying the number of nodes $n\in\{80,100,120,140\}$ and the degree $d\in\{2,5,10\}$. Specifically, for each pair $(n,d)$, we generate $50$ instances. Thus, in total, we have $50\times 4\times 3=600$ instances in the test set. If there is no further explanation about the hardware setting, the hardware graph is set in form of Chimera topology which is a $45\times 45$ grid of $K_{4,4}$ unit cells (Boothby et al., 2016). The experiment on the synthetic dataset aims to evaluate the performance of CHARME on logical graphs of the same distribution.
Furthermore, we also conducted the experiment on a real dataset generated based on the QUBO formulations of test instances of the Target Identification by Enzymes (TIE) problem proposed in (Ngo et al., 2023b). In particular, different metabolic networks for the TIE problem result in different QUBO formulations. As we mentioned in previous sections, each specific QUBO formulation corresponds to a logical graph. Thus, we can generate various logical graphs by varying the metabolic networks for the TIE problem. In this experiment, we collect $185$ metabolic networks from the KEGG database (Kanehisa and Goto, 2000). Then, we generate the corresponding logical graphs using the QUBO formulations with these metabolic networks as input. In the collected graphs, the number of nodes ranges from $6$ to $165$, while the number of edges ranges from $22$ to $578$. We split the dataset into three groups based on the ratio between the number of nodes $n$, and the number of edges $m$: low density group ($m\leq 2n$), medium density group ($2n<m\leq 3n$), and high density group ($m>3n$). We use $70$ graphs for training and preserve the remaining for testing. We denote the training set $\mathcal{G}_{real}$. This experiment aimed to illustrate the effectiveness of our hybrid RL model in real-world scenarios.
Benchmark. To demonstrate the effectiveness of our Order Exploration strategy, we establish a comparative greedy strategy referred to as Greedy Exploration. In Greedy Exploration, nodes in the logical graph are selected for embedding in a recursive manner at each step. The order of node selection for recursion is determined by the number of additional qubits required to embed each node. Additionally, recursion at a given step is terminated if the number of embedded qubits exceeds the best solution found up to that point.
@@
  • OCT-based (Goodrich et al., 2018): Following the top-down approach, it minimizes the number of qubits based on their proposed concept of the virtual hardware framework. In this experiment, we use the most efficient version of OCT-based, named Fast-OCT-reduce. Although OCT-based is effective in reducing the number of qubits, it suffers from high computational complexity, leading to a long running time to find a solution. This can pose a bottleneck for the QA process, especially when the frequency of QA usage has increased rapidly.
  • Minorminer (Cai et al., 2014): A heuristic approach developed commercially by D-Wave, this method finds solutions by consecutively adding nodes of the logical graph to a chain in the hardware graph without considering the isolated problem (a node in the hardware graph can be embedded by more than one node from the logical graph). Then, the solution with overlapping embedded nodes is refined by iteratively removing and adding better chains. The algorithm terminates when the time limit is reached or a feasible solution is achieved. Minorminer is an efficient approach for sparse logical graphs.
@@
  • ATOM (Ngo et al., 2023a): This approach is a heuristic method that finds solutions based on the concept of adaptive topology. The strength of this method is the ability to find solutions in short running times without compromising their quality. In addition, ATOM uses a heuristic strategy to decide the order of embedding nodes. Thus, by comparing its performance to that of ATOM, we can assess the effectiveness of RL in leveraging the order selection strategy.
  • ATOM (random): This method is a variant of ATOM presented in (Ngo et al., 2023a). The difference is that ATOM (random) uses a random strategy to decide the embedding order. Specifically, the next node of the logical graph selected to be embedded in the hardware graph is chosen randomly from the set of unembedded nodes of the logical graph.
  • ATOM (degree): This method is another variant of ATOM presented in (Ngo et al., 2023a). The difference is that ATOM (degree) uses a degree-based strategy to decide the embedding order. Specifically, the next node of the logical graph selected to be embedded is the node with the highest degree from the set of unembedded nodes of the logical graph.
@@
Hyper-parameter Selection. In the exploration phase, we apply Order Exploration, described in Algorithm 2, to precompute embedding orders for each graph $P_{i}\in\mathcal{G}_{syn}\cup\mathcal{G}_{real}$. The constants $\beta_{i}$ for $i\in[1,m]$ are determined as the solutions for $P_{i}$ found by ATOM (Ngo et al., 2023a). Specifically, we set the sampling limit $D$ as $10000$, the exploration limit $K$ as $10^{6}$ ,and use up to $1024$ threads running in parallel to explore embedding orders.
@@
For training, we set the number of $GCN$ layers and fully connected layers for both actor and critic models as $L=3$ and $C=3$, respectively. The learning rate for the actor model is $3\times 10^{-4}$, while the learning rate for the critic model is $10^{-5}$. The number of episodes to update the policy is $100$, and the batch size for each update is $290$.
@@
For inference, due to the stochasticity of ML models, we sample $10$ solutions for each test case, and report the solution with the least number of qubits. In addition, we use multiprocessing to sample solutions, so the actual running time of inference is equal to the time to sample one solution.
@@
Figure 4. The figure highlights the performance of Order Exploration and Greedy Exploration in discovering embedding orders for training graphs in synthetic training sets under various settings of $(n,d)$. In subfigures, the x-axis denotes the number of exploration steps, while the y-axis represents the corresponding efficiency scores of the resulting orders. As a result, in these subfigures, each orange (blue) point indicates the best efficiency score of the embedding order resulted by Order Exploration (Greedy Exploration) after a specific number of exploration steps.
Figure 5. The figure highlights the performance of Order Exploration and Greedy Exploration in discovering embedding orders for training graphs in the real training sets under various graph types including low density, medium density, and high density. In subfigures, the x-axis denotes the number of exploration steps, while the y-axis represents the corresponding efficiency scores of the resulting orders. As a result, in these subfigures, each orange (blue) point indicates the best efficiency score of the embedding order resulted by Order Exploration (Greedy Exploration) after a specific number of exploration steps.
4.2. Results
4.2.1. The performance of the proposed exploration strategy
This section demonstrates the utility of Order Exploration in generating precomputed actions, commonly referred to as embedding orders, to streamline the training process of CHARME. Our experiments are designed to assess two key aspects: 1) the effectiveness of Theorem 1 with a general prefix that implies the practical significance of the lower bound outlined in Equation 17, and 2) the overall performance of our proposed exploration strategy in terms of the efficiency scores of the resulting embedding orders.
@@ 4.2.1. The performance of the proposed exploration strategy
The Effectiveness of Theorem 1 with a General Prefix. In this experiment, we designed a meticulous procedure to evaluate the effectiveness of Theorem 1 for general prefix orders, which are not necessarily non-expansion orders. Specifically, we compare the number of qubits gained when embedding the same node $v$ at two different embedding steps $t$ and $t+\delta$, denoted as $C(v|t)$ and $C(v|t+\delta)$. In Theorem 1, we prove that $C(v|t)\leq C(v|t+\delta)$ if there is no expansion embedding between $t$ and $t+\delta$. Thus, our focus here lies on comparing $C(v|t)$ and $C(v|t+\delta)$ when there is at least one expansion embedding between these steps. Given a graph $P=(V_{P},E_{P})\in\mathcal{G}_{syn}\cup\mathcal{G}_{real}$, the procedure for selecting $v$, $t$, and $\delta$ unfolds as follows. First, we randomly select a subset $S_{P}\subset V_{P}$ by selecting a random prefix $O_{A}\in\mathcal{E}^{(S_{P})}$. We select $v$ as a random node in $V_{P}\setminus S_{P}$. Next, we need to select $t$ and $\delta$ such that there exists an expansion embedding from $t$ to $t+\delta$. In detail, we start by iteratively embedding nodes following the prefix order $O_{A}$. Subsequently, we randomly embed nodes $v^{\prime}\in V_{P}\setminus S_{P},v^{\prime}\neq v$ until we obtain an expansion embedding at step $|S_{P}|+\tau$. Finally, $t$ is assigned as $|S_{P}|+\tau-1$ and $\delta$ is randomly selected such that $\delta\geq 1$. Performing this procedure repeatedly, we can obtain different $v$, $t$, and $\delta$ for each training graph. For each triplet $(v,t,\delta)$ of a given training graph, $C(v|t)$ and $C(v|t+\delta)$ is calculated similarly as discussed in Theorem 1. For each setting of $(n,d)$, we measure the average of $C(v|t)$ and $C(v|t+\delta)$ on all triplet $(v,t,\delta)$ sampled from graphs with the setting $(n,d)$. We denote the average on $C(v|t)$ and $C(v|t+\delta)$ as early embedded gain and late embedded gain, respectively.
@@
The results are illustrated in Table 2. Across all $(n,d)$ configurations, the average qubit requirement for embedding a node prior to topology expansion (early embedded gain) is consistently lower compared to post-topology expansion (late embedded gain). Interestingly, when the graph density increases, the difference in qubit gain between before and after topology expansion becomes more significant. This trend can be attributed to the expansion of the topology, which, on average, elongates the length of clean paths. Consequently, embedding a node after topology expansion tends to incur higher costs. These observations highlight the significance of the lower bound $\bar{F}$ on the efficiency score of embedding orders with a general prefix, as derived from Equation 17.
@@
The Efficiency in Exploring Orders. In this experiment, we conducted a comparison between Order Exploration and Greedy Exploration to assess their effectiveness in exploring embedding orders for training logical graphs across various node sizes and degrees $(n,d)$. Figure 4 presents subfigures corresponding to each setting $(n,d)$ of the logical graphs in the synthetic training set $\mathcal{G}_{syn}$. Each subfigure displays the efficiency scores of the embedding orders resulted from both strategies with variations in the exploration limit $K$ which is mentioned in Algorithm 2. We recall that given a logical graph $P$ and a hardware graph $H$, the efficiency score of an embedding order $O$ is the number of qubits when embedding $P$ into $H$ following the order $O$.
@@
The results depicted in the subfigures highlight the superior performance of Order Exploration over Greedy Exploration across all $(n,d)$ settings. In detail, the disparity becomes more significant with increasing values of $n$ and $d$. For example, with $(n,d)=(80,2)$, the average efficiency score of Order Exploration is approximately $2\%$ less than that of Greedy Exploration, while with $(n,d)=(140,10)$, the gap increases to around $9\%$. Furthermore, we observe that the disparity also widens as the exploration limit $K$ increases from 0 to $10^{6}$.
The primary difference between the two methods lies in their recursion termination conditions. Specifically, Order Exploration utilizes a termination condition based on the lower bound on embedding non-expansion prefixes, as derived from Theorem 1. In contrast, Greedy Exploration does not incorporate this lower bound in its termination condition. Consequently, Theorem 1 proves to be practically useful for guiding the exploration of embedding orders, particularly in scenarios involving large $(n,d)$ values and high exploration limits $K$.
@@
Similarly, when applied to the real training set $\mathcal{G}_{real}$, the Order Exploration approach consistently outperforms Greedy Exploration in terms of efficiency scores (Figure 5). Specifically, when dealing with low-density and medium-density graphs, the disparity between Order Exploration and Greedy Exploration is small. This is because the limited number of nodes and edges restricts the variation in qubit usage across different embedding orders. For example, in order to embed logical graphs with fewer nodes and edges, the number of required expansion steps, which consumes significantly more qubits than non-expansion steps, is small. Thus, the qubit usage between different embedding orders shows little variation. However, when dealing with high-density graphs, Order Exploration significantly outperforms Greedy Exploration. Specifically, Order Exploration requires approximately $1100$ qubits, while Greedy Exploration requires more than $1600$ qubits for its solution. This illustrates the superior performance of Order Exploration, particularly in settings involving high-density graphs.
4.2.2. The limitation of NaiveRL
Here, we evaluate the practical performance of NaiveRL to identify its limitation. In this experiment, we adopt a different setting. Specifically, we generate a dataset of $1000$ graphs and split them into training and test sets. Our goal is to train the model on $700$ BA graphs with $n=70$ nodes and degree $d=5$, aiming for effective generalization to the remaining $300$ graphs, which have varying sizes ($n\in\{30,50\}$) and a lower degree ($d=2$).
We employ the Actor-Critic method (Mnih et al., 2016) to update the policy, following Equations 6 and 7. Both the actor and critic are implemented as fully connected neural networks. During training, graphs from the training set are sequentially selected as initial states for exploration. The exploration history is then recorded and used to update the policy after all graphs in the training set have been processed.
Figure 6 highlights the limitations of NaiveRL in producing feasible solutions. Without the exploration term in Equation 8, NaiveRL struggles to generate feasible solutions during training, with a feasibility rate of less than $10\%$. Incorporating the exploration term improves feasibility by encouraging the model to mimic precomputed solutions, leading to a higher feasibility rate during training. However, this improvement does not generalize well to unseen graphs, where the feasibility rate drops to below $50\%$. Notably, NaiveRL encounters difficulties even with very small graphs. This suggests that its performance will deteriorate further as graphs become denser and larger. Thus, a naive RL approach is inadequate for addressing the minor embedding problem in quantum annealing, which fundamentally requires feasible solutions for every input graph.
To address these challenges, we propose CHARME, an enhanced RL framework designed to generalize effectively to unseen graphs while guaranteeing the feasibility of its solutions. In the following section, we demonstrate the effectiveness of CHARME through comprehensive evaluations.
@@ 4.2.3. The Performance of CHARME
In this experiment, we compare the performance of CHARME against three state-of-the-art methods on synthetic and real datasets.
Figure 7. Training performance in term of average qubit usage of CHARME with and without precomputed solutions from Order Exploration.
n 80 100 120 140 degree 2 5 10 2 5 10 2 5 10 2 5 10 OCT-based 357.76 845.38 1222.11 515.74 1263.32 1850.9 725.84 1782.18 2607.06 931.24 2348.74 3488.18 Minorminer 338.56 1236.88 2110.16 501.56 1914.18 3313.24 664.76 2632.82 4854.10 830.56 3554.01 6573.84 ATOM 384.06 1126.14 1815.72 546.62 1782.86 2882.72 727.46 2560.36 4217.48 998.58 3448.24 5621.88 ATOM (random) 456.90 1411.14 2304.69 683.44 2146.00 3713.32 932.74 3089.38 5398.51 1234.72 4179.40 7476.58 ATOM (degree) 549.64 1482.06 2481.76 840.84 2242.20 3845.93 1096.64 3109.12 5530.90 1407.96 4103.83 7313.83 CHARME 324.62 1027.36 1686.20 485.16 1580.08 2566.14 655.88 2263.66 3870.02 821.88 3082.59 5100.54
@@
Table 3. The average qubit usage of OCT-based, Minorminer, ATOM, and CHARME corresponding to 12 different graph sizes in the synthetic dataset.
@@
Figure 8. Comparison between four methods in terms of running time in the synthetic dataset.
Figure 9. Comparison between four methods in terms of success rate with different hardware size limits. Hardware used in this experiment is in the form of $C(N,N,4)$ (an $N\times N$ grid of bipartite units $K_{4,4}$)
@@
The Synthetic Dataset. Here, we assess the performance of CHARME on the synthetic dataset, focusing on the number of qubits used and the running time. The goals of this experiment are as follows:
  • We examine the training performance of CHARME with and without the precomputed solutions resulting from Order Exploration.
@@
  • We evaluate how effectively CHARME leverages the heuristic strategies to decide the order of the node embedding used in ATOM and its variants.
@@
  • We compare the quality of the solutions in terms of the number of qubits obtained by CHARME with three state-of-the-art methods.
@@
  • We measure the inference time of CHARME and compare it with the running time of other methods to assess its computational efficiency.
  • We assess the success rate for finding embedding solutions in a short running time with different hardware size limits of CHARME and other methods.
Figure 7 illustrates the training performance of CHARME in terms of average qubit usage in two scenarios: without and with precomputed solutions from Order Exploration. In the ”without precomputed solutions” scenario, we remove the exploration term on the reward function. First, the results demonstrate that precomputed solutions accelerate CHARME’s exploration for high-quality solutions. Specifically, without support from precomputed solutions, CHARME gets stuck on random solutions (i.e., average qubit usage $\approx 8200$) in the first $12000$ steps before eventually discovering good patterns (with average qubit usage dropping to $7200$). On the other hand, when using precomputed solutions, CHARME merges quickly to those solutions within the first $1000$ steps (i.e., average qubit usage $\approx 7800$). With a better starting point, CHARME explores high-quality patterns more rapidly, with the average qubit usage dropping to $6900$ by the $9000$th step. In addition, we can see how efficient CHARME can explore in the solution space. In detail, after $30000$ steps of training, CHARME reduces the number of qubits required to embed graphs in the training set by approximately $25\%$.
@@
Table 3 presents the average number of qubits obtained from solutions generated by four methods for various pairs of size and degree $(n,d)$ of logical graphs in the test set. First, the results imply that CHARME leverages the heuristic strategies for the order of node embedding used in ATOM and its variants. Across all pairs $(n,d)$, CHARME consistently provides solutions with fewer qubits compared to ATOM. In particular, CHARME achieves reductions in qubit usage of up to $11.36\%$, $68.21\%$, and $72.78\%$ compared to ATOM, ATOM (random), and ATOM (degree), respectively. This observation indicates the advantage of using RL for determining the order of node embedding compared to heuristic and naive strategies. Additionally, the learning ability of the proposed RL framework is evident, as the policy learned from the training set performs well on new graphs of different sizes and degrees.
@@
Turning to solution quality, the results in Table 3 demonstrates that CHARME outperforms three state-of-the-art methods in terms of qubit usage in test cases with degree $d=2$. On average, in these cases, CHARME achieves $2.52\%$ reduction in terms of qubit usage, compared to the best current methods. In other cases, although CHARME does not overcome the OCT-based approach in terms of qubit usage, it still outperforms Minorminer and ATOM. This observation, combined with the subsequent analysis of the running time, is important to illustrate the practical effectiveness of CHARME in handling the minor embedding problem.
@@
Next, we examine the running time of OCT-based, Minorminer, ATOM, and CHARME. We note that the running time of ATOM, ATOM (random), and ATOM (degree) is identical, so we do not show the running time of ATOM (random) and ATOM (degree) in this experiment. The results in Figure 8 illustrate that the running time of OCT-based is significantly larger than the running time of three other methods. In particular, the running time of OCT-based is approximately 6 times bigger than the running time of the second slowest method, Minorminer. Thus, although OCT-based is efficient in terms of qubit usage in some certain cases, its expensive running time can pose a significant bottleneck in practical QA processes. On the other hand, the running time of CHARME is comparable to the fastest method, ATOM. Furthermore, as mentioned above, CHARME can offer better solutions than Minorminer and ATOM in all cases. Therefore, CHARME can be the optimal choice for tackling the minor embedding problem in practice, especially when current QA processes must handle user requests immediately due to increased demand for their use.
Finally, we assess the success rate in finding feasible solutions of fast embedding methods, including CHARME, Minorminer, ATOM, ATOM (random), and ATOM (degree) under different hardware size limits. The hardware used in this experiment is in the form of an $NxN$ grid of $K_{4,4}$ unit cells with $N\in\{5,10,15,20,25,30,35\}$. The running time limit in this experiment is set as $10$ seconds. The success rate is defined as the proportion of test cases (out of 600) in which a feasible solution is found within the running time limit. We note that Minorminer includes a post-processing function that aims to refine the final solution to fit the hardware size constraint, while CHARME, ATOM, ATOM (random), and ATOM (degree) do not include any function to refine solutions. The results in Figure 9 illustrate that CHARME achieves a significantly higher success rate than ATOM, ATOM (random), and ATOM (degree) in all seven cases. In the case of $N=10$, the success rate of CHARME is twice that of ATOM and three times that of ATOM (random). This demonstrates the superiority of RL over heuristic and naive strategies in determining the order of node embedding by significantly boosting the success rate. Furthermore, CHARME’s success rate is comparable to that of Minorminer (even achieves a higher success rate in two cases of $N=25$ and $N=30$). This observation indicates the efficiency and potential of CHARME in finding feasible solutions with limited hardware resources. Specifically, by determining a decent order of node embedding, CHARME achieves a success rate comparable to that of Minorminer without requiring refinement of the final solution.